# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v2

@@ V Information theoretical aspects
What we have noted in the immediately preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of and- and sum-definitions. However, what decreases in case of mixed 2ESS( or-definition) is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$). In passing, we may note that an even stricter condition on KL-divergence is compatible with and-definition: It is obvious that the fixed point corresponding to ESS of and-definition, leads to monotonic decrease of $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})$ and $D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ separately.
@@
We would like to make a passing remark here that a well-celebrated concept in the theory of statistical inference, viz., Fisher information Fisher (1922) metric (${\sf J}$)—which represents a metric of distance between two nearby probability distributions—is related to the Shahshahani metric: In symmetric games, by mapping the probability mass function $(p(1),p(2),\cdots,p(n))\to(x_{1},x_{2},\cdots,x_{n})$, we can easily show Harper (2009) that the Fisher information metric gets converted into Shahshahani metric:
@@
Our aim is to search for a potential $U$ such that the following is true (a potential whose Shahshahani gradient is replicator equation):
@@
since $\bm{\xi}$ and $\bm{\eta}$ being tangent vectors, satisfy $\sum_{i}\xi_{i}=0$ and $\sum_{i}\eta_{i}=0$, respectively. As far as the R.H.S. of Eq. (54) is concerned, one can easily verify that following is the potential whose Shahshahani gradient is replicator dynamics by substituting it in the R.H.S. of Eq. (53):
@@
To arrive at this conclusion, we must invoke the properties of ${\sf\Omega}$ and ${\sf\Omega}^{\prime}$ that $\sum_{i}\xi_{i}({\sf\Omega}\mathbf{y})_{i}=0$ and $\sum_{i}\eta_{i}({\sf\Omega}^{\prime}\mathbf{x})_{i}=0$ owing to the respective facts that $({\sf\Omega}\mathbf{y})_{i}$ and $({\sf\Omega}^{\prime}\mathbf{x})_{i}$ are independent of the index $i$.
In the next step, we construct a Lyapunov function $V$ from the potential $U$. Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$, if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0~{}\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $V$ takes the minimum value—in fact, 00. Now, extrema of $V$ is attained at a point where
@@
This means that the extrema $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is nothing but the internal fixed point of the replicator dynamics, Eq. (34). For this fixed point to be minima, the Hessian matrix,
@@
We can see that ${\sf H}$ is a symmetric matrix which if positive definite makes the extremum at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ a minima Kuttler (2008). Supposing a minimum exists at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, we evaluate the time derivative of $V$ to be
@@
for all the neighborhood ($\mathbf{x},\mathbf{y}$) of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$). Now replacing ($\mathbf{x}$, $\mathbf{y}$) by $(2\mathbf{x}-\mathbf{\hat{x}},2\mathbf{y}-\mathbf{\hat{y}}$), which is also in the neighbourhood Hofbauer and Sigmund (1998) of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$), we get,
@@ VI ESS in Multiplayer hypermatrix games
In the preceding sections, we delved into the definitions of ESS and elucidated their respective domains of applicability, i.e., we clarified which definitions are useful for characterizing pure ESS and which are useful for mixed ESS, all within the context of two-player interactions. However, what if the strategic interaction involves more than two players? In such scenarios, we must shift our focus to the framework of multiplayer games Broom et al. (1997); Broom and Rychtář (2022); Gokhale and Traulsen (2010). Within the realm of multiplayer games, questions might arise, what is the definition of ESS? Are there multiple definitions of ESS, and what are their respective domains of applicability? In this section, we address these questions. Multiplayer games can be somewhat intricate. For the purpose of this paper, it suffices to focus—unless otherwise specified—on multiplayer hypermatrix games where the payoffs are linear in their respective arguments Broom and Rychtář (2022).
In multiplayer games, the payoff hypermatrix depends on multiple arguments. We denote ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ as the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies. If $d_{i}$ opponents are of same strategy, $\mathbf{p}_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of $\mathbf{p}_{i}$-strategied opponent.
@@
Introducing a short hand notation for average population strategy as $\tilde{\mathbf{p}}=\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}}$, inequality (63) can be rewritten as, ${\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$. (Note that $\tilde{\mathbf{p}}$ is in the infinitesimal neighbourhood of $\mathbf{\hat{p}}$, for sufficiently small $\epsilon$.) We multiply both sides of inequality (63) by $\epsilon$ and add $(1-\epsilon){\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$ to both sides to get,
@@
Absorbing the $\epsilon$ and $(1-\epsilon)$ inside the arguments of the payoff, we get
@@
However, note that those definitions are in terms of the states and not strategies. When the above formalism of hypermatrix games is transferred to a population with phenotypes identified with pure strategies, say $\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}$, the fitness of $i$-type may be denoted as ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$, where $k$-th component ($k=1,2,\cdots,m$) of the $m$-dimensional vector $\mathbf{x}$ represents the frequency of $k$-th strategy. Here, we have adopted a notational convention: When $\mathbf{p}_{i}$ appears inside square brackets, as in ${\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]$, it represents the unit vector in the direction of the $i$-th pure state on the simplex $\Sigma_{m}$, rather than the strategy $\mathbf{p}_{i}$ itself. The value of ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$ is given by $\sum_{{i_{2}}{i_{3}},\cdots,{i_{\nu}}}x_{i_{2}}x_{i_{3}},\cdots,x_{i_{\nu}}{\sf U}\left[\mathbf{p}_{i};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$. Similarly, ${\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\equiv\sum_{{i_{1}}{i_{2}},\cdots,{i_{\nu}}}x_{i_{1}}x_{i_{2}},\cdots,x_{i_{\nu}}{\sf U}\left[\mathbf{p}_{i_{1}};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$, and ${\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]\equiv\sum_{{i_{1}}{i_{2}},\cdots,{i_{\nu}}}\hat{x}_{i_{1}}x_{i_{2}},\cdots,x_{i_{\nu}}{\sf U}\left[{\mathbf{p}}_{i_{1}};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$. Finally, in the context of states we can say that a state $\mathbf{\hat{x}}$ is an Evolutionary Stable State (ESS) if
@@
To explore asymptotic stability of its fixed point, we propose the following Lyapunov function: $V_{s}(\mathbf{x})=\sum_{i}\hat{x}_{i}\log({\hat{x}_{i}}/{x_{i}})$. This function is positive definite by definition of KL-divergence and $V_{s}(\mathbf{\hat{x}})=0$. We evaluate
@@ VI.1 Asymmetric hypermatrix games
  1. Definition 1e (and-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{and}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (72)
@@
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{or}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (74)
@@ VI.1.1 Pure ESS
This inequality has to hold for all values of $(\mathbf{x},\mathbf{y})$ in some neighbourhood of the ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$. If the ESS is not pure, then we can choose $x_{2}=\hat{x}_{2}$, $\cdots$, $x_{n-1}=\hat{x}_{n-1}$ and $y_{2}=\hat{y}_{2}$, $\cdots$, $y_{n-1}=\hat{y}_{n-1}$. Thus, one can always set the unfixed components $x_{1}$ and $y_{1}$ in such a manner that sign of the remaining term can be flipped; this, however, is not possible if the ESS is pure as $\hat{x}_{1}$ and $\hat{y}_{1}$ can be either zero or one. This proves the assertion. In fact, same assertion is valid for any ESS under sum-definition (see Appendix B).
Of course, in general asymmetric hypermatrix games, ESS’s under all three definitions can be either pure or mixed. The next logical question is to ask ESS (of either bihypermatrix game or asymmetric hypermatrix game) under which definition is (locally) asymptotically stable fixed point of the replicator equation (71). As far as the pure ESS’s under and- and sum-definitions are concerned, the same Lyapunov function as used in the case of two-player game (see Sec. III.2), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq. (71). Thus, one can show that and- and sum-definitions imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is take the time derivative of $V_{1}$ and use Eq. (71) to arrive at
@@
and subsequently demand the $\dot{V}_{1}<0$ to validate the implication. Moreover, following exactly similar arguments as done in the two-player case (Sec. IV.1), linear stability analysis establishes that pure 2ESS is asymptotically stable fixed point of the replicator equation (71) even in the multiplayer case when it is a strict NE. Using the same argument we can also show that all three definitions are equivalent in case of strict NE. We omit the straightforward steps as they are analogous and repetitive, as one can see from the eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are— $({\sf A}[\mathbf{p}_{1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf A}[\mathbf{p}_{2};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{2};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),\cdots,({\sf A}[\mathbf{p}_{m-1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf C}[\mathbf{q}_{1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{1};\mathbf{q}_{n}^{{\gamma_{y}}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),({\sf C}[\mathbf{q}_{2};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{2};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),\cdots,({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{n-1};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}])$, at the fixed point $(\hat{\mathbf{x}},\hat{\mathbf{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$.

### v3

@@ V Information theoretical aspects
What we have noted in the preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of and- and sum-definitions. However, what decreases in case of mixed 2ESS( or-definition) is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$). In passing, we may note that an even stricter condition on KL-divergence is compatible with and-definition: It is obvious that the fixed point corresponding to ESS of and-definition, leads to monotonic decrease of $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})$ and $D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ separately.
@@
We would like to make a passing remark here that a well-celebrated concept in the theory of statistical inference, viz., Fisher information Fisher (1922) metric (${\sf J}$)—which represents a metric of distance between two nearby probability distributions—is related to the Shahshahani metric: In symmetric games, by mapping the probability mass function $(p(1),p(2),\cdots,p(n))$ to $(x_{1},x_{2},\cdots,x_{n})$, we can easily show Harper (2009) that the Fisher information metric gets converted into Shahshahani metric:
@@
Our aim is to search for a potential $U$ such that the following is true (a potential whose Shahshahani gradient is the replicator equation):
@@
since $\bm{\xi}$ and $\bm{\eta}$, being tangent vectors, satisfy $\sum_{i}\xi_{i}=0$ and $\sum_{i}\eta_{i}=0$, respectively. As far as the R.H.S. of Eq. (51) is concerned, one can easily verify that it is the R.H.S of Eq. (50) with the function:
@@
So $U$ is the potential whose Shahshahani gradient is the replicator dynamics. To arrive at this conclusion, we must invoke the properties of ${\sf\Omega}$ and ${\sf\Omega}^{\prime}$ that $\sum_{i}\xi_{i}({\sf\Omega}\mathbf{y})_{i}=0$ and $\sum_{i}\eta_{i}({\sf\Omega}^{\prime}\mathbf{x})_{i}=0$ owing to the respective facts that $({\sf\Omega}\mathbf{y})_{i}$ and $({\sf\Omega}^{\prime}\mathbf{x})_{i}$ are independent of the index $i$.
In the next step, we construct a Lyapunov function $V$ from the potential $U$. Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$. If $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0\leavevmode\nobreak\ \forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $V$ takes the minimum value which is 00. Now an extremum of $V$ is attained at a point where
@@
This means that the extremum $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is nothing but the internal fixed point of the replicator dynamics (31). For this fixed point to be minimum, the Hessian matrix,
@@
We can see that ${\sf H}$ is a symmetric matrix which if positive definite makes the extremum at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ a minimum Kuttler (2008). Supposing a minimum exists at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, we evaluate the time derivative of $V$ to be
@@
for every $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ in an appropriate neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$. Now replacing ($\mathbf{x}$, $\mathbf{y}$) by $(2\mathbf{x}-\mathbf{\hat{x}},2\mathbf{y}-\mathbf{\hat{y}}$), which is also in the neighbourhood of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) (see Ref. Hofbauer and Sigmund (1998)), we get,
@@ VI ESS in Multiplayer hypermatrix games
In the preceding sections, we delved into the definitions of ESS and elucidated their respective domains of applicability, i.e., we clarified which definitions are useful for characterizing pure ESS and which are useful for mixed ESS, all within the context of two-player interactions. However, what if the strategic interaction involves more than two players? In such scenarios, we must shift our focus to the framework of multiplayer games Broom et al. (1997); Broom and Rychtář (2022); Gokhale and Traulsen (2010). Within the realm of multiplayer games, the questions might arise what is the definition of ESS? Are there multiple definitions of ESS, and what are their respective domains of applicability? In this section, we address these questions. Multiplayer games can be somewhat intricate. For the purpose of this paper, it suffices to focus—unless otherwise specified—on multiplayer hypermatrix games where the payoffs are linear in their respective arguments Broom and Rychtář (2022).
In multiplayer games, the payoff hypermatrix depends on multiple arguments. We denote by ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies. If $d_{i}$ opponents are of same strategy, $\mathbf{p}_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of an opponent employing strategy $\mathbf{p}_{i}$.
@@
Introducing a short hand notation for average population strategy as $\tilde{\mathbf{p}}=\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}}$, inequality (60) can be rewritten as, ${\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$. (Note that $\tilde{\mathbf{p}}$ is in the infinitesimal neighbourhood of $\mathbf{\hat{p}}$, for sufficiently small $\epsilon$.) We multiply both sides of inequality (60) by $\epsilon$ and add $(1-\epsilon){\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]$ to both sides to get,
@@
We mention that as done in two-player symmetric games, one can choose an $\bar{\epsilon}$ independent of $\mathbf{p}$ such that inequality (60) is valid for all $\epsilon<\bar{\epsilon}$, since the union of all faces of the simplex that exclude $\mathbf{\hat{p}}$ is compact (see Ref. Broom and Rychtář (2022); Bukowski and Miekisz (2004)). Now, absorbing the $\epsilon$ and $(1-\epsilon)$ inside the arguments of the payoff, we get
@@
However, note that those definitions are in terms of the states and not strategies. When the above formalism of hypermatrix games is transferred to a population with phenotypes identified with pure strategies, say $\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}$, the fitness of $i$-type may be denoted as ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$, where the $k$-th component ($k=1,2,\cdots,m$) of the $m$-dimensional vector $\mathbf{x}$ represents the frequency of the $k$-th strategy. Here, we have adopted a notational convention: When $\mathbf{p}_{i}$ appears inside square brackets, as in ${\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]$, it represents the unit vector in the direction of the $i$-th pure state on the simplex $\Sigma_{m}$, rather than the strategy $\mathbf{p}_{i}$ itself. The value of ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$ is given by $\sum_{{i_{2}}{i_{3}},\cdots,{i_{\nu}}}x_{i_{2}}x_{i_{3}},\cdots,x_{i_{\nu}}{\sf U}\left[\mathbf{p}_{i};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$. Similarly, ${\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\equiv\sum_{{i_{1}}{i_{2}},\cdots,{i_{\nu}}}x_{i_{1}}x_{i_{2}},\cdots,x_{i_{\nu}}{\sf U}\left[\mathbf{p}_{i_{1}};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$, and ${\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]\equiv\sum_{{i_{1}}{i_{2}},\cdots,{i_{\nu}}}\hat{x}_{i_{1}}x_{i_{2}},\cdots,x_{i_{\nu}}{\sf U}\left[{\mathbf{p}}_{i_{1}};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$. Finally, in the context of states we can say that a state $\mathbf{\hat{x}}$ is an Evolutionary Stable State (ESS) if
@@
To explore asymptotic stability of its fixed points, we propose the following Lyapunov function: $V_{s}(\mathbf{x})=\sum_{i}\hat{x}_{i}\log({\hat{x}_{i}}/{x_{i}})$. This function is positive definite by definition of KL-divergence and $V_{s}(\mathbf{\hat{x}})=0$. We evaluate
@@ VI.1 Asymmetric hypermatrix games
  1. Definition 1e (and-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ {\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (69)
@@
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]\leavevmode\nobreak\ \text{\emph{or}}\leavevmode\nobreak\ {\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (71)
@@ VI.1.1 Pure ESS
This inequality has to hold for all values of $(\mathbf{x},\mathbf{y})$ in some neighbourhood of the ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$. If the ESS is not pure, then we can choose $x_{2}=\hat{x}_{2}$, $\cdots$, $x_{n-1}=\hat{x}_{n-1}$ and $y_{2}=\hat{y}_{2}$, $\cdots$, $y_{n-1}=\hat{y}_{n-1}$. Thus, one can always set the unfixed components $x_{1}$ and $y_{1}$ in such a manner that sign of the remaining term can be flipped; this, however, is not possible if the ESS is pure as $\hat{x}_{1}$ and $\hat{y}_{1}$ can be either zero or one. This proves the assertion. In fact, same assertion is valid for any ESS under sum-definition (see Appendix C).
Of course, in general asymmetric hypermatrix games, ESSes under all three definitions can be either pure or mixed. The next logical question is to ask ESS (of either bihypermatrix game or asymmetric hypermatrix game) under which definition is (locally) asymptotically stable fixed point of the replicator equation (68). As far as the pure ESSes under and- and sum-definitions are concerned, the same Lyapunov function as used in the case of two-player game (see Sec. III.2), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq. (68). Thus, one can show that and- and sum-definitions imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is to take the time derivative of $V_{1}$ and use Eq. (68) to arrive at
@@
and subsequently demand the $\dot{V}_{1}<0$ to validate the implication. Moreover, following exactly similar arguments as done in the two-player case (Sec. IV.1), linear stability analysis establishes that pure 2ESS is asymptotically stable fixed point of the replicator equation (68) even in the multiplayer case when it is a strict NE. We omit the straightforward steps as they are analogous and repetitive, as one can see from the eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are— $({\sf A}[\mathbf{p}_{1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf A}[\mathbf{p}_{2};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{2};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),\cdots,({\sf A}[\mathbf{p}_{m-1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf C}[\mathbf{q}_{1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{1};\mathbf{q}_{n}^{{\gamma_{y}}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),({\sf C}[\mathbf{q}_{2};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{2};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),\cdots,({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{n-1};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}])$, at the fixed point $(\hat{\mathbf{x}},\hat{\mathbf{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$. Using the same argument we can also show that all three definitions are equivalent in case of strict NE.